# Enhanced Droplet Analysis Using Generative Adversarial Networks

### v1

@@ Abstract
Precision devices play an important role in enhancing production quality and productivity in agricultural systems. Therefore, the optimization of these devices is essential in precision agriculture. Recently, with the advancements of deep learning, there have been several studies aiming to harness its capabilities for improving spray system performance. However, the effectiveness of these methods heavily depends on the size of the training dataset, which is expensive and time-consuming to collect. To address the challenge of insufficient training samples, this paper proposes an alternative solution by generating artificial images of droplets using generative adversarial networks (GAN). The GAN model is trained by using a small dataset captured by a high-speed camera and capable of generating images with progressively increasing resolution. The results demonstrate that the model can generate high-quality images with the size of $1024\times 1024$. Furthermore, this research leverages recent advancements in computer vision and deep learning to develop a light droplet detector using the synthetic dataset. As a result, the detection model achieves a 16.06% increase in mean average precision (mAP) when utilizing the synthetic dataset. To the best of our knowledge, this work stands as the first to employ a generative model for augmenting droplet detection. Its significance lies not only in optimizing nozzle design for constructing efficient spray systems but also in addressing the common challenge of insufficient data in various precision agriculture tasks. This work offers a critical contribution to conserving resources while striving for optimal and sustainable agricultural practices.
Keywords: Polyps, Colonoscopy, Medical image analysis, Deep learning, Vision transformers.
@@ I-A Motivation
Precision spraying systems play a crucial role in modern agriculture due to their impact on droplet drift and efficacy. Therefore, to enhance precision in spraying systems, it’s vital to optimize spray nozzles. This involves adjusting nozzle parameters to change the characteristics of released droplets and make them suitable to each crop’s unique needs and environmental conditions. There are several methods to measure the droplet characteristics. Traditional optical measurement methods include high-speed photography and laser imaging [1, 2]. For example, one common method estimates droplet size by spraying colored water on a white sheet and then analyzing the resultant patterns [3, 4]. These methods are often highly manual, require delicate setups, and are associated with high operating costs.
Recently, with the advent of advanced machine learning, [5] developed a novel tool to estimate the size and velocity of droplets sprayed from agricultural nozzles. Moreover, [6] developed a lightweight droplet detector that can be easily incorporated into sensing systems. Despite these advancements, these models are heavily dependent on the size of the datasets used for model training. However, gathering data to train machine learning models in most precision agriculture applications remains a time-consuming, labor-intensive, and expensive process.
@@ I-B Literature review
Over the past decade, deep learning has been developing rapidly with appealing applications across various domains, including finance, medicine, education, and agriculture [7, 8, 9, 10]. In agriculture, these techniques have played a crucial role in addressing complex challenges such as crop detection, fruit segmentation, and produce classification. For instance, [11] applied deep-learning algorithms for the detection and segmentation of apples. For crop detection, [12] developed an algorithm capable of tracking crops in sunny conditions.
Beyond the application of machine learning in crop treatment, it also plays an important role in obtaining measurements for optimizing the design of agricultural devices. For example, work in [5, 13] employed deep-learning techniques to estimate the spray cone angles and the characteristics of the droplets. To use these novel methods, data collection is crucial and always a challenge because it requires complicated and expensive set-up and excessive labor costs. Therefore, the accuracy and reliability of the model often depend on the training dataset in terms of both quality and quantity.
To resolve this data-size bottleneck and hence increase the performance of deep learning models, there are several techniques we can use, for example collecting more data, transfer learning, data augmentation, and synthetic data generation [14]. For data augmentation, we can apply transformations to existing data to create new samples, including rotations, flips, zooms, or changes in brightness and contrast. Generally, the performance of the model could be improved by applying the data augmentation techniques. Compared to the data augmentation techniques, synthetic data generation can create new samples based on the original samples. This can be especially beneficial when creating novel patterns or dealing with data from class imbalance.
Generative Adversarial Networks (GANs) have revolutionized content and model generation by introducing a novel framework for training generative models [15]. A GAN model is usually constructed by two models: a generator and a discriminator. The generator produces images from noise to fool the discriminator, while the discriminator aims to distinguish real images from generated ones. This mechanism is trained until the images generated by the generator resemble real images.
@@
The application of generative models in agriculture has become an alternative to traditional methods, leading to a substantial enhancement in precision agriculture [23, 24, 25]. For example, work in [26] created a new dataset of artificial plants for plant phenotyping by using a generative model. Furthermore, the study in [25] demonstrates that they can achieve an accuracy of 95% using only synthetic data for crop seed phenotyping. By using GAN models, time and resources can be saved compared to traditional data collection methods, which are the key inspirations for this paper.
@@ I-C Contributions
In this study, to address the bottleneck of data limitation in our previous work [5], we propose a method to generate synthetic droplet images via developing a GAN model. Since droplets are small, generating small objects that have a similar appearance as droplets poses a tremendous challenge for a generative model. Therefore, our approach involves generating high-resolution images from low-resolution ones by leveraging the progressive generative methodology. Additionally, this paper establishes a novel droplet detection method using a cutting-edge deep-learning algorithm trained on the AI-generated synthetic data to measure their size.
@@
  • To the best of our knowledge, this study is the first attempt to leverage deep learning techniques for generating droplet images and establishing a substantial synthetic droplet dataset.
  • The droplet image generator in this work is capable of producing images with a resolution of $1024\times 1024$, which is higher than the original images.
@@ III-A1 Generative mechanism
$\begin{split}u_{i}=\min_{G}\max_{D}([&\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log(D(x))]+\\ &\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))])].\end{split}$
@@ IV-A Evaluation metrics for GAN
To assess the quality of synthetic images produced by the generative model, we employ a qualitative assessment method, and this method contains two essential properties: Fidelity and Diversity. Fidelity is used for assessing the quality of the generated samples, measuring how realistic the images appear. Diversity evaluates the range of the generated samples, assessing how well they cover the entirety of the real distribution’s variety [30].
The fidelity and diversity of the synthesized images generated by the generator are evaluated by precision and recall metrics through droplet detection. High precision indicates that the generated images closely resemble real images on average, while high recall suggests that the generator can produce any sample present in the training dataset [31].
@@ IV-B Evaluation metrics for Object detection
$\text{AP}=\int_{0}^{1}\text{Precision}(\text{Recall})\,d(\text{Recall})$ (4)
@@ V-A Synthetic Dataset with the generative model
Figure 7: Droplet images produced by the generative model.
As we observed, these images exhibit a notable resemblance to real ones, illustrating not only in terms of resolution but also in capturing intricate details, including small droplets. The generative model demonstrates its capability to produce realistic droplet images, thereby enriching the database with a diverse array of visual features.
@@ VI Conclusion
This work introduces a novel approach to streamline spray system optimization through a data augmentation method. The method uses a generative adversarial neural network to create a synthetic dataset. The generative model consists of a generator and a discriminator: the generator produces images from noise to fool the discriminator; the discriminator tries to distinguish images generated by the generator and the real images.
To train the generative AI model, we collect data from experiments using a high-speed camera to capture droplet images. Instead of initially training the generator to produce high-resolution images, the training process begins with the generator creating images at a small resolution of $4\times 4$ and gradually increasing it up to $1024\times 1024$. This progressive training facilitates the model in learning features at different scales, enhancing its stability and ability to generate high-quality images.

### v3

@@ Abstract
Precision devices play an important role in enhancing production quality and productivity in agricultural systems. Therefore, the optimization of these devices is essential in precision agriculture. Recently, with the advancements of deep learning, there have been several studies aiming to harness its capabilities for improving spray system performance. However, the effectiveness of these methods heavily depends on the size of the training dataset, which is expensive and time-consuming to collect. To address the challenge of insufficient training samples, we developed an image generator named DropletGAN to generate images of droplets. The DropletGAN model is trained by using a small dataset captured by a high-speed camera and capable of generating images with progressively increasing resolution. The results demonstrate that the model can generate high-quality images with the size of $1024\times 1024$. The generated images from the DropletGAN are evaluated using the Fréchet inception distance (FID) with an FID score of 11.29. Furthermore, this research leverages recent advancements in computer vision and deep learning to develop a light droplet detector using the synthetic dataset. As a result, the detection model achieves a 16.06% increase in mean average precision (mAP) when utilizing the synthetic dataset. To the best of our knowledge, this work stands as the first to employ a generative model for augmenting droplet detection. Its significance lies not only in optimizing nozzle design for constructing efficient spray systems but also in addressing the common challenge of insufficient data in various precision agriculture tasks. This work offers a critical contribution to conserving resources while striving for optimal and sustainable agricultural practices.
Keywords: Droplet detection, Deep learning, Computer vision, GAN, Precision agriculture, Sprayer systems.
@@ I-A Motivation
Precision spraying systems play a crucial role in modern agriculture due to their impact on droplet drift and efficacy [1, 2]. There are several kinds of sprayers such as air blast sprayers [3], steam-powered sprayers [4], smart sprayers [5], etc. The air blast sprayer, for instance, comes in various configurations, including axial-fan, cannon, tunnel, tower, and custom designs. While axial-fan sprayers, with their large radial spray volume, have been the mainstay for tree fruit orchards, they have limitations in modern, intensive settings. With axial-fan sprayers, they are susceptible to significant off-target drift and chemical waste due to their broad spray pattern [1].
Smart sprayers, on the other hand, are a new generation of sprayers that leverage advanced technologies like machine learning and computer vision to improve the efficiency and precision of chemical applications in agriculture [6]. Unlike traditional sprayers that apply a uniform amount of chemicals across a field, smart sprayers can adjust the application rate based on real-time data. This data can include information about the presence and type of weeds, the size and health of crops, and even weather conditions [7]. By targeting only the areas that need treatment, smart sprayers can significantly reduce waste and environmental impact, while also minimizing costs for farmers. Additionally, some smart sprayers can detect and target individual plants, further reducing chemical use and exposure for workers [8].
Optimizing spray nozzles is crucial for enhancing precision in spraying systems. This optimization involves adjusting nozzle parameters to change the characteristics of released droplets and make them suitable to each crop’s unique needs and environmental conditions [9]. There are several methods to measure the droplet characteristics. Traditional optical measurement methods include high-speed photography and laser imaging [10, 11]. For example, one common method estimates droplet size by spraying colored water on a white sheet and then analyzing the resultant patterns [12, 13]. These methods are often highly manual, require delicate setups, and are associated with high operating costs.
Recently, with the advent of advanced machine learning, [14] developed a novel tool to estimate the size and velocity of droplets sprayed from nozzles. However, the methods are not suitable for measuring the droplet deposition on stems and leaves in field spraying. Moreover, [15] developed a lightweight droplet detector that can be easily incorporated into sensing systems. [16] proposed a droplet deposition characteristics detection method based on deep learning. Despite these advancements, these models are heavily dependent on the size of the datasets used for model training. However, gathering data to train machine learning models in most precision agriculture applications remains a time-consuming, labor-intensive, and expensive process [17, 18, 6].
@@ I-B Literature review
Over the past decade, deep learning has been developing rapidly with appealing applications across various domains, including finance, medicine, education, and agriculture [19, 20, 21, 22, 23]. In agriculture, deep learning techniques have played a crucial role in addressing complex challenges such as crop detection, fruit segmentation, and produce classification [24, 25, 26]. For instance, [27] applied deep-learning algorithms for the detection and segmentation of apples. For crop detection, [28] developed an algorithm capable of tracking crops in sunny conditions. For weed detection, [29] proposed a real-time detection method using deep learning. For harvesting, [30] proposed a deep learning-based object detection for detecting coconut clusters.
Beyond the application of machine learning in crop treatment, it also plays an important role in obtaining measurements for optimizing the design of agricultural devices. For example, work in [14, 31] employed deep-learning techniques to estimate the spray cone angles and the characteristics of the droplets. Moreover, [32] proposed a deep-learning approach for discerning the droplet jetting dynamics in single-jet printing processes. However, a major challenge in utilizing these novel methods lies in data collection. The process often requires complex and expensive setups, incurring significant labor costs. As a result, the accuracy and reliability of the resulting machine learning models heavily depend on the quality and quantity of data within the training dataset.
To resolve this data-size bottleneck and hence increase the performance of deep learning models, there are several techniques we can use, for example collecting more data, transfer learning, data augmentation, and synthetic data generation [33]. For data augmentation, we can apply transformations to existing data to create new samples, including rotations, flips, zooms, or changes in brightness and contrast [33]. Generally, the performance of the model could be improved by applying the data augmentation techniques. Compared to the data augmentation techniques, synthetic data generation can create new samples based on the original samples. This can be especially beneficial when creating novel patterns or dealing with data from class imbalance.
Generative Adversarial Networks (GANs) have revolutionized content and model generation by introducing a novel framework for training generative models [34]. GANs are usually constructed by two models: a generator and a discriminator. The generator produces images from noise to fool the discriminator, while the discriminator aims to distinguish real images from generated ones. This mechanism is trained until the images generated by the generator resemble real images.
@@
In addition to GANs, Variational Autoencoder (VAEs) and Flow-based Deep Generative Models are commonly used for image generation [42, 43]. For example, in VAEs, the models compress an input into a latent space representation, capturing its essence, and then attempt to reconstruct the original data from this compressed version. In VAEs, the twist lies in how the models represent the data in latent space as a probability distribution. This probabilistic approach allows VAEs to not only reconstruct data but also to sample new points from the learned distribution, effectively generating novel data that resembles the training data. Despite its efficiencies and performances, VAEs can sometimes struggle with capturing complex data structures.
In 2020, the field of diffusion models saw a significant advancement with the proposal of denoising diffusion probabilistic models (DDPM) by [44]. Prior to this, several generative models based on the core concept of diffusion had been introduced, such as those by [45, 46]. In diffusion models, we start by gradually adding noise to a real image from the training data. This process essentially corrupts the image with increasing levels of noise across a series of steps. Then, we train a diffusion model to learn the inverse process, which is how to remove the added noise and recover the original clean image gradually. Once trained, these models can generate entirely new images by starting with pure noise and progressively removing noise following the learned denoising process. With each step, the noise is reduced, revealing more and more image details. Diffusion models distinguish themselves from GANs and VAEs by utilizing a fixed training procedure and employing high-dimensional latent variables.
Diffusion models have emerged as a powerful tool for generating different kinds of data. For image generation, DALL-E 2, developed by OpenAI, is one the popular models that create high-fidelity images from text descriptions [47]. Similar to DALL-E 2, Imagen is another powerful text-to-image diffusion model developed by Google AI [48]. In addition, Stable Diffusion is also built using denoising diffusion probabilistic techniques [49]. Beyond image generation, active research is exploring diffusion models’ capabilities in diverse areas such as generating videos, training time improvement, speech generation, and image reconstruction [50, 51, 52, 53].
The application of generative models in agriculture has become an alternative to traditional methods, leading to a substantial enhancement in precision agriculture [54, 55, 56]. For example, [57] created a new dataset of artificial plants for plant phenotyping by using a generative model. [56] demonstrates that they can achieve an accuracy of 95% using only synthetic data for crop seed phenotyping. [58] proposed an augmentation technique using diffusion models for weed recognition.
Despite its performance, diffusion-based models require a large dataset, and it takes a long time to train, especially with complex structured datasets. While VAEs and flow-based models offer alternative approaches, they can struggle to capture complex details and generate high-resolution images. Therefore, we proposed a droplet generator using GAN architectures, named DropletGAN. Our approach aims to save time and resources compared to traditional data collection methods for droplets.
@@ I-C Contributions
In this study, to address the bottleneck of data limitation and droplet analysis, we propose DropletGAN, a method used to generate synthetic droplet images. Since droplets are small, generating small objects that have a similar appearance as droplets poses a tremendous challenge for a generative model. Therefore, our approach involves generating high-resolution images from low-resolution ones by leveraging the progressive generative methodology. Additionally, this paper establishes a novel droplet detection method using a cutting-edge deep-learning algorithm trained on AI-generated synthetic data to measure their size.
@@
  • We developed DropletGAN, which is capable of generating high-quality droplet images.
  • The Fréchet inception distance (FID) score of the generated images is 11.29.
@@ III-A1 Generative mechanism
$\begin{split}u_{i}=\min_{G}\max_{D}([&\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log(D(x))]+\\ &\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))])].\end{split}$ (1)
@@ IV-A Evaluation metrics for GAN
To assess the quality of synthetic images produced by the generative model, we use the Fréchet inception distance (FID) [62]. FID compares the distributions of features extracted from real images and those generated by our GAN. A lower FID score suggests the generated images are more realistic and diverse, closely resembling the distribution of real images. In FID, a pre-trained Inception V3 model is used to extract features from both real images (denoted by X) and generated images (denoted by G) [63]. This process essentially encodes the images into a lower-dimensional vector space capturing their characteristics. Then we model the data distribution for these features using a multivariate Gaussian distribution with mean $\mu$ and covariance $\Sigma$. The FID is computed using the following formula:
[…]
Here, $||\mu_{X}-\mu_{G}||^{2}$ represents the squared Euclidean distance between the mean vectors ($\mu_{X}$ and $\mu_{G}$) of the real and generated features. $\text{Tr}(\Sigma_{X}+\Sigma_{G}-2(\Sigma_{X}^{\frac{1}{2}}\Sigma_{G}^{\frac{1}{2}}))$ represents the trace of the sum of the covariance matrices ($\Sigma_{X}$ and $\Sigma_{G}$) of the real and generated data, minus twice the product of their square roots. $\Sigma_{X}^{\frac{1}{2}}$ denotes the matrix square root of the covariance matrix for the real data. $\Sigma_{G}^{\frac{1}{2}}$ denotes the matrix square root of the covariance matrix for the generated data.
In addition, the fidelity and diversity of the synthesized images generated by the generator are evaluated by precision and recall metrics through droplet detection. High precision indicates that the generated images closely resemble real images on average, while high recall suggests that the generator can produce any sample present in the training dataset [64, 65].
@@ IV-B Evaluation metrics for Object detection
$[h]\text{AP}=\int_{0}^{1}\text{Precision}(\text{Recall})\,d(\text{Recall})$ (6)
@@ V Results and discussion
Figure 7: Droplet images produced by the generative model.
@@ V-A Synthetic Dataset with the generative model
The calculated FID score is 11.29, which is small enough to generate good-quality images [66, 67]. From observation, the generated images exhibit a notable resemblance to real ones, illustrating not only in terms of resolution but also in capturing intricate details, including small droplets. The generative model demonstrates its capability to produce realistic droplet images, thereby enriching the database with a diverse array of visual features.
@@ VI Conclusion
This work introduces a novel approach to streamline spray system optimization through a data augmentation method. We proposed DropletGAN to create a synthetic dataset. The DropletGAN model consists of a generator and a discriminator: the generator produces images from noise to fool the discriminator; the discriminator tries to distinguish images generated by the generator and the real images.
To train the generative AI model, we collect data from experiments using a high-speed camera to capture droplet images. Instead of initially training the generator to produce high-resolution images, the training process begins with the generator creating images at a small resolution of $4\times 4$ and gradually increasing it up to $1024\times 1024$. This progressive training facilitates the model in learning features at different scales, enhancing its stability and ability to generate high-quality images. As a result, the Fréchet inception distance score of the generated dataset is 11.29. This score illustrates that these images are good-quality images compared to real images.